OpenStudy (anonymous):

If line PQ = 3 units, PR = 8 units and PT = 12 units. how long is PS?

OpenStudy (anonymous):



OpenStudy (anonymous):

here u have to use PR*PQ =PT*PS

OpenStudy (anonymous):

thus PS=PR*PQ/PT

OpenStudy (anonymous):

PS = 8x3 / 12 24/12 PS=2units Did I get it right?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 8 }{ 3 }=\frac{ 12 }{ x }\] \[x=\frac{ 12*3 }{ 8 }=\frac{ 36 }{ 8 }=4.5\]

OpenStudy (anonymous):

Can I ask what theorem or postulate is being used?

OpenStudy (anonymous):

i just placed corresponding sides equal

OpenStudy (anonymous):

i mean their ratios

OpenStudy (anonymous):

Okay then, how long is Pr if Pu = 12cm and PQ = 6cm ?

OpenStudy (anonymous):

\[\frac{ 12 }{ PR }=\frac{ 12 }{ 4.5}\] PR=4.5

OpenStudy (anonymous):

why did you use 4.5 ?

OpenStudy (anonymous):

@Roshee 2 is correct for the question u posted

OpenStudy (anonymous):

Was that the Cross-Products Property?

OpenStudy (anonymous):

@Roshee how long is Pr if Pu = 12cm and PQ = 6cm in this case use PQ.PR =PU*PU and hence PR= PU*PU/PQ

OpenStudy (anonymous):

why multiply 2 PU ?

OpenStudy (anonymous):

PR = 12 x 12 divided by 6cm = 144 / 6 = 24cm